A simple building with walls of unknown construction has been determined to have a super insulated roof, which allows essentially no heat transfer. You are to determine the heat transfer coefficient for the walls. The walls measure 1 foot thick, and the building is of rectangular shape 20 feet on the long side in 10 feet on the short side with ceilings 8 feet high. From the heating bills for the month of January, in which the average outside temperature was 30°F, it was determined that the house loses heat at the rate of 5000 BTU/hour. During that month the interior temperature was held at an average of 68°F. Since the building has only two small windows and one door the effect of these on heat transfer can be neglected for the purposes of this initial study
Answer #1

s Heat transter through walls = 5000 BTU/hr 20 Heat transferred through Walls = _KADT I A = QA1+QA2 = 2x8x20 + 2x8x10 A = 480